# Chain-of-Frames: Advancing Video Understanding in Multimodal LLMs via Frame-Aware Reasoning

### v1

@@ 4.1 Experimental setup
Models selection. While chain-of-frames is a general approach, we find it particularly well-suited for the recent InternVL models (Chen et al., 2024; Zhu et al., 2025). These models introduce a novel format for videos, where frames are interleaved with text identifiers such as Frame-1, Frame-2, etc. (see Fig. 5), reinforcing the temporal structure of the video. This format already associates the image with their corresponding textual identifiers, and using these identifiers in reasoning traces may facilitate long-range interaction in long-context LLMs. The InternVL models achieve state-of-the-art performance among open-weight video LLMs, and thus improving their reasoning capabilities is both challenging and of practical relevance. Consequently, we adopt the recently released InternVL2.5-4B (Chen et al., 2024) and InternVL3-8B (Zhu et al., 2025) models for our experiments. We leave exploration of CoF on other architectures to future work.
@@
Training details. For InternVL2.5-4B, we fully fine-tune both the LLM and projection modules, while keeping the vision encoder frozen. For InternVL3-8B, we apply LoRA-based fine-tuning (Hu et al., 2022) to reduce memory usage. Detailed training configurations can be found in App. A.
@@
Video benchmarks. We compare the video LLMs on five popular benchmarks that capture diverse aspects of video understanding. These benchmarks span a broad range of tasks, video types and duration, providing a comprehensive evaluation of model capabilities. Video-MME (Fu et al., 2024) includes six visual domains, including videos from 2 to 60 minutes long. MVBench (Li et al., 2023b) encompasses 20 tasks that require more than single-frame analysis. VSI-Bench (Yang et al., 2024) focuses on quantitative reasoning tasks such as object counting or appearance order. To evaluate hallucination tendencies of video LLMs, we additionally include the VidHal (Choong et al., 2024) and EventHallusion (Zhang et al., 2024a) benchmarks. Details on each benchmark and the breakdown of the results over the fine-grained splits are available in App. A and App. B.
@@
Original + CoT Prompting 33.5 54.7 71.5 77.0 67.4
SFT with QA only 31.8 54.5 73.4 64.1 57.7
SFT with CoT 34.3 58.6 73.7 77.9 53.1
SFT with CoF (ours) 36.9 59.7 76.1 79.2 71.2
4.2 Chain-of-Frames vs other Chain-of-Thoughts variants
@@
  • Original: the InternVL2.5-4B model with default prompting.
  • Original + CoT Prompting: the InternVL2.5-4B model with a prompt that encourages the model to perform intermediate reasoning before answering the question (see prompt in App. A). This approach is similar to the standard CoT used in language tasks (Wei et al., 2022).
  • SFT with QA only: the InternVL2.5-4B model fine-tuned on the question-answers pairs from our CoF-Data without including the reasoning traces.
  • SFT with CoT: the InternVL2.5-4B model fine-tuned on CoF-Data, where reasoning traces are included but references to specific frames are removed (e.g., “In Frame 1…” is replaced with a generic “In the video…”). This approach mimics the standard CoT format and is similar to the method of Fei et al. (2024). As their model is not publicly available, this method provides a direct comparison to their approach (VideoCoT) on our more comprehensive dataset.
  • SFT with CoF: the InternVL2.5-4B model fine-tuned on CoF-Data, i.e. our proposed approach.
@@
Results. We report results for all models on the five benchmarks in Table 1. First, we observe that CoT prompting alone already improves the accuracy of the original model on four out of five benchmarks compared to standard prompting. This result highlights the value of encouraging explicit reasoning, which video LLMs are capable of performing. Second, fine-tuning on question-answer pairs without reasoning (SFT with QA only) gives mix results, possibly due to overfitting on the training data, which might degraded the reasoning ability of the original InternVL2.5-4B. Training on reasoning traces without temporal grounding (SFT with CoT) improves the results across nearly all benchmarks, with the notable exception of EventHallusion. Finally, the model trained on the full CoF-Data (SFT with CoF), i.e., fine-tuned on the reasoning traces including frame references, achieves the highest accuracy across all benchmarks, with improvements ranging from 4.8% to 8.7% over the original InternVL2.5-4B. These results demonstrate that our chain-of-frames approach, which combines reasoning with temporal grounding, offers consistent benefits across diverse tasks and domains, from quantitative reasoning to hallucination detection. Finally, as shown in App. C, the reasoning with frame references provide interpretable insights into the model’s decision-making process.
4.3 Comparison to state-of-the-art video LLMs
We want now to contextualize the performance of our CoF-based models with the current leading video LLMs. For this, in Table 2 we report the performance of some of the strongest both closed-source (GPT-4V/4T (Achiam et al., 2023), GPT-4o (Hurst et al., 2024), Gemini-1.5-Pro (Gemini Team et al., 2024)) and open-source (Qwen2-VL-72B (Wang et al., 2024a), LLaVA-OneVision-72B (Li et al., 2024a)) models of different size. We could not add the original models from other works (Wang et al., 2024b; Han et al., 2024; Fei et al., 2024; Hu et al., 2025b) as these are not publicly available. However, we provide comparisons based on the common benchmarks in App. B, where our CoF-models outperform these baselines. Besides the InternVL2.5-4B fine-tuned on CoF-Data, we provide an additional CoF-based LLMs obtained by fine-tuning the more recent InternVL3-8B: we indicate them as CoF-InternVL2.5-4B and CoF-InternVL3-8B respectively. We note that we do not refer to the results of any models on our benchmarks as zero-shot. In fact, for closed-source and some open-source models, the training data is not known. Moreover, 5 out 20 tasks of MVBench are partially based on the test set of Clevrer and VidHal includes some videos from MVBench, and, e.g., InternVL2.5 models have been pre-trained on Clevrer.
Table 2 shows that, despite being significantly smaller than most baselines, CoF-InternVL3-8B achieves the best results on VSI-Bench, Video-MME, and MVBench, outperforming leading (closed-source) video LLMs. Moreover, it has second-best accuracy on VidHal, and best among open-source LLMs on EventHallusion. Meanwhile, CoF-InternVL2.5-4B attains competitive results, surpassing several larger open-source models and in some cases even closed-source ones. These results demonstrate the effectiveness of using chain-of-frames reasoning to improve performance across diverse and challenging video understanding benchmarks. Finally, in App. B we report the breakdown of the results over the subsets of the benchmarks in Table 2. From those we can see, for example, that training with CoF leads to improvements in the spatial reasoning tasks of VSI-Bench. We hypothesize that this is due to having the reasoning traces about object count, appearance order, and relative distance created from Clevrer in CoF-Data. This suggests that the model is effectively learning spatial reasoning tasks from synthetic CoF data.
Table 2: Comparison of CoF-models to state-of-the-art video LLMs. We report accuracy on the five evaluation benchmarks for relevant baseline models and our CoF-models fine-tuned on CoF-Data. Both CoF-InternVL2.5-4B and CoF-InternVL3-8B outperform the majority of baselines despite significantly fewer parameters, and obtain the best performance on three out of five benchmarks.
Model VSI-Bench Video-MME MVBench VidHal EventHall
Closed-source Models
GPT-4V/4T - 59.9 43.7 - 76.5
GPT-4o 34.0 71.9 - 77.2 91.9
Gemini-1.5-Pro 48.8 75.0 - 67.1 80.4
Open-source Models
LLaVA-OneVision-72B 40.2 66.2 59.4 64.7 59.5
Qwen2-VL-72B 37.6 71.2 73.6 76.2 54.7
LLaVA-OneVision-7B 32.4 58.2 56.7 58.4 60.1
LLaVA-NeXT-Video-7B 35.6 46.5 53.1 50.9 64.8
Qwen2-VL-7B 31.0 63.3 67.0 69.6 59.3
InternVL2.5-4B 33.5 54.7 71.5 77.0 67.4
InternVL3-8B 41.0 66.5 74.4 80.9 72.1
Our Models
CoF-InternVL2.5-4B 36.9 59.7 76.1 79.2 71.2
CoF-InternVL3-8B 51.3 75.3 90.4 79.5 78.7
4.4 Additional analyses of Chain-of-Frames
Figure 6: Diversity of training data. Combining CoF examples extracted from real and synthetic videos provides better results on almost all benchmarks compared to using a single source. For all models we use 164k training samples, and fine-tune an InternVL2.5-4B model.
Influence of training data.
The training data of CoF-Data combines chain-of-frames from real and synthetic videos (total of 164k samples). To evaluate the impact of dataset diversity, we construct datasets of the same size (164k samples) but from a single source, i.e., either CoF-Datareal or CoF-Datasynth, and fine-tune InternVL2.5-4B on them. We compare these two models to standard CoF-InternVL2.5-4B (trained on CoF-Data) in Fig. 6. Using the combined dataset outperforms single-source datasets on all benchmarks except for EventHallusion, demonstrating the importance of diversity in the reasoning traces used for training. Between the two individual datasets, the model trained on synthetic videos outperforms the one trained on real videos in three out of five cases, while being worse in just one. This suggests that further improvements might come from expanding the tasks covered by the synthetic data, which can even be generated at low cost.
@@ 4.4 Additional analyses of Chain-of-Frames
CoF reasoning at inference time. To better understand how the CoF models utilize frame references during inference, we track the number of frames referenced in the reasoning trace per answer. In Fig. 7, we report how many answers, generated by our CoF-InternVL2.5-4B and CoF-InternVL3-8B models across all evaluation benchmarks, contain $N$ frames for $N=1,\ldots,10$. We can see that both models learn to produce, from a relatively small training set, reasoning traces which refer to frame IDs (see qualitative examples in App. C). In particular, CoF-InternVL2.5-4B includes at least one frame reference in 68.7% of the answers, while CoF-InternVL3-8B does it in 76.9% of the cases. This suggests that larger models might benefit more from fine-tuning with frame-grounded reasoning: in fact, the CoF-InternVL3-8B model shows both higher use of frame references and larger gains in accuracy (see Fig. 1). Finally, the models can use frame references selectively rather than uniformly across tasks ( distribution by benchmark in App. B), depending on the type of questions.
@@
Figure 7: CoF reasoning at inference time. For both CoF-InternVL2.5-4B (left plot) and CoF-InternVL3-8B (right), we show the distribution of the number of frame references generated at inference time: both models learn to produce chain-of-frames during evaluation.
@@ 5 Conclusion
We have introduced chain-on-frames (CoF), a new approach to encourage video LLMs to produce temporally grounded reasoning before providing the final answer. Compared to existing works, CoF does not require complex ad-hoc inference frameworks or auxiliary models, and we show that its training data can be efficiently extracted from both real and synthetic videos. Our models fine-tuned on CoF-Data outperform across multiple benchmarks those obtained with alternative methods for reasoning, and even achieve results better than or similar to leading video LLMs. Overall, these features make CoF a viable option to further improve the reasoning capabilities of video LLMs. Exploring the effect of increasing the size and diversity of the training data, as well as the scale of the models, represents an exciting direction for future work.
Limitations. We have currently applied CoF on the InternVL models since their video encoding is well-suited for our reasoning with frame references, and they achieve state-of-the-art performance. It is an interesting open question how to customize our approach to other types of video LLMs.
@@ A.1 Chain-of-Frames training data
CoF from real videos (CoF-Datareal). To generate question-reasoning-answer triplets, we prompt Llama-3.1-8B-Instruct (Meta AI, 2024) using the instruction shown in Fig. 8 along with frame-aware video captions from the VideoEspresso dataset (see Fig. 3 for details). Notably, the raw video content is not included in this process. Two examples from CoF-Datareal are shown in Fig. 12.
CoF from synthetic videos (CoF-Datasynth). The second portion of our training dataset is derived from the Clevrer dataset, which includes detailed attributes for each object in every video frame. Specifically, given a frame ID and object ID, the inside_camera_key field indicates whether the object is visible in the frame, enabling us to determine when an object enters or exits the scene. The velocity attribute reflects whether an object is moving or stationary, while the location attribute provides its absolute or relative position, which can be leveraged to estimate distances or identify collisions. The final CoF-Datasynth dataset comprises three categories of questions: Object Count, appearance order, and relative distance. Within the object count category, we define three subtypes: (i) collision-based (“How many collisions…”), (ii) motion state (“How many moving objects…”), and (iii) temporal-based, where questions reference specific segments of the video (“After object A enters…”). The questions, answers and reasoning traces are generated with the manually written templates shown in Fig. 9, making the data collection process particularly simple and fast. Examples from CoF-Datasynth are shown in Fig. 13.

### v2

@@ 4.1 Experimental setup
Model VSI-Bench Video-MME MVBench VidHal EventHall Average
Closed-source Models
GPT-4V/4T - 59.9 43.7 - 76.5 -
GPT-4o 34.0 71.9 - 77.2 91.9 -
Gemini-1.5-Pro 48.8 75.0 - 67.1 80.4 -
Open-source Models
LLaVA-OneVision-72B 40.2 66.2 59.4 64.7 59.5 58.0
Qwen2-VL-72B 37.6 71.2 73.6 76.2 54.7 62.7
LLaVA-OneVision-7B 32.4 58.2 56.7 58.4 60.1 53.2
LLaVA-NeXT-Video-7B 35.6 46.5 53.1 50.9 64.8 50.2
Qwen2-VL-7B 31.0 63.3 67.0 69.6 59.3 58.0
InternVL2.5-4B 33.5 54.7 71.5 77.0 67.4 60.8
InternVL3-8B 41.0 66.5 74.4 80.9 72.1 67.0
Our Models
CoF-InternVL2.5-4B 36.9 59.7 76.1 79.2 71.2 64.6
CoF-InternVL3-8B 51.3 73.7 77.1 79.5 78.7 72.1
Table 1: Comparison of CoF-models to state-of-the-art video LLMs. We report accuracy on five benchmarks (and their average) for relevant baselines and our CoF-models fine-tuned on CoF-Data. Both CoF-InternVL2.5-4B and CoF-InternVL3-8B outperform the majority of baselines despite significantly fewer parameters, and obtain the best or second-best results on four out of five benchmarks.
@@
Baseline models. While chain-of-frames is a general approach, we find it particularly well-suited for the recent InternVL models [6, 43]. These models introduce a novel format for videos, where frames are interleaved with text identifiers such as Frame-1, Frame-2, etc. (see Fig. 4), reinforcing the temporal structure of the video. This format already associates the image with their corresponding textual identifiers, and using these identifiers in reasoning traces may facilitate long-range interaction in long-context LLMs. The InternVL models achieve state-of-the-art performance among open-weight video LLMs, and thus improving their reasoning capabilities is both challenging and of practical relevance. Consequently, we adopt the recently released InternVL2.5-4B [6] and InternVL3-8B [43] models for our experiments. To test the generalizability of our approach beyond the InternVL family, we use we employ Phi-3.5-Vision [1], a mobile-scale multimodal LLM that demonstrates strong performance in language reasoning, as a third baseline model. For limited spaces, we present the results relative to Phi-3.5-Vision in App. B, where we show that CoF is also effective on this video LLM.
@@
Training details. For InternVL2.5-4B, we fully fine-tune both the LLM and projection modules, while keeping the vision encoder frozen. For InternVL3-8B, we apply LoRA-based fine-tuning [15] to reduce memory usage. We refer to the resulting fine-tuned models as CoF-InternVL2.5-4B and CoF-InternVL3-8B throughout the paper. Detailed training configurations can be found in App. A.
Inference details. We uniformly sample 30 frames from each input video for inference. This approach ensures consistent temporal coverage regardless of video length. Importantly, the model is not restricted to 30-second videos and can process videos of arbitrary duration.
@@
Video benchmarks. We compare the video LLMs on five popular benchmarks that span a broad range of tasks, video types and duration, providing a comprehensive evaluation of model capabilities. Video-MME [10] includes six visual domains, including videos from 2 to 60 minutes long. MVBench [26] encompasses 20 tasks that require more than single-frame analysis. VSI-Bench [38] focuses on quantitative reasoning tasks like object counting or appearance order. To evaluate hallucination in video LLMs, we include VidHal [8] and EventHallusion [40]. Details on each benchmark and the breakdown of the results over the fine-grained splits are available in App. A and App. B.
4.2 Main experiments
We now contextualize the performance of our CoF-based models in relation to the current leading video LLMs. For this, in Tab. 1 we report the performance of some of the strongest both closed-source (GPT-4V/4T [2], GPT-4o [19], Gemini-1.5-Pro [11]) and open-source (Qwen2-VL-72B [31], LLaVA-OneVision-72B [23]) models of different sizes. We could not add the prior video CoT models [32, 14, 9, 17] in Tab. 1 as these are not publicly available. Nevertheless, Tab. 3 shows comparisons on shared benchmark datasets, where our CoF-based models outperform these baselines. Finally, we note that we do not refer to the results of any models on our benchmarks as zero-shot. In fact, for closed-source and some open-source models, the training data is not known. Moreover, 5 out 20 tasks of MVBench are partially based on the test set of Clevrer and VidHal includes some videos from MVBench.
Comparison to base models. Tab. 1 shows that CoF-enhanced models consistently outperform their corresponding base models across nearly all benchmarks. CoF-InternVL3-8B achieves 5.1% higher average performance than the original InternVL3-8B, while the CoF-InternVL2.5-4B improves of 3.8% over InternVL2.5-4B. Moreover, the fact that performance gains are more notable for the larger model suggests that more capable LLMs may benefit more substantially from CoF-based fine-tuning.
Comparison to state-of-the-art video LLMs. Tab. 1 also illustrates that, despite being significantly smaller than most baselines, CoF-InternVL3-8B achieves the best results on VSI-Bench and MVBench, outperforming leading (closed-source) video LLMs. It also has second-best accuracy on Video-MME and VidHal, and best among open-source LLMs on EventHallusion. Meanwhile, CoF-InternVL2.5-4B attains competitive results, surpassing several larger open-source models and in some cases even closed-source ones. These results demonstrate the effectiveness of chain-of-frames reasoning in improving video understanding across diverse and challenging benchmarks.
Comparison to video CoT methods. Prior video CoT models [32, 14, 9, 17] are not accessible to the public, and they do not report results on the five benchmarks we use. However, we provide comparative analyses to M-LLM [17] and Video-of-Thought [9] on Video-MME and NextQA [36] using the results they report. Both baselines adopt multi-stage inference that explicitly retrieves key frames, whereas our approach integrates frame retrieval implicitly through frame references within the reasoning traces. In Tab. 3, we see that our CoF-based models, including CoF-InternVL2.5-4B, which has fewer parameters than the competitors, outperform these baselines on both benchmarks. Moreover, compared to the most recent model, M-LLM, our method achieves a greater improvement (4.9% vs 0.8%) on NextQA despite starting from a stronger base model.
5 Additional analyses
5.1 Influence of synthetic training data
Figure 6: Influence of synthetic training data. Across most benchmarks, the model trained on synthetic data consistently outperforms the one trained on real data. Despite the substantial distribution gap between the synthetic samples and the real frames in these benchmarks, the model effectively learns the tasks from synthetic data and transfers this knowledge to real-world scenarios. For all models we use 164k training samples, and fine-tune an InternVL2.5-4B model.
The training data of CoF-Data combines chain-of-frames from real and synthetic videos. As mentioned in Sec. 3.3, the synthetic portion of the dataset, CoF-Datasynth, is generated without relying on heavy models, but rather through human-designed templates. The key question is how effective this synthetic data is for model training.
To address this question, we construct datasets of equal size, each sourced exclusively from either CoF-Datareal or CoF-Datasynth, and fine-tune InternVL2.5-4B on them. We then compare the resulting CoF-InternVL2.5-4B models in Fig. 6. First, we observe that across most benchmarks, the model trained on synthetic data consistently outperforms the one trained on real data. This result is notable: despite the considerable distributional gap between the synthetic samples and real frames from the benchmarks, the model effectively learns tasks such as object counting and appearance ordering from synthetic examples. Since synthetic data can be generated at low cost and large scale, this finding highlights a promising and scalable strategy for improving performance. Second, using the combined dataset outperforms single-source datasets on all benchmarks except for EventHallusion, demonstrating the importance of diversity in the reasoning traces used for training.
5.2 Chain-of-frames vs chain-of-thoughts variants
Backbone Model Video-MME NextQA
Video-LLaVA-7B Original 39.9 66.3
[…]
Qwen2-VL-7B Original 58.1 77.6
[…]
InternVL2.5-4B Original 54.9 75.3
[…]
InternVL3-8B Original 66.5 82.4
[…]
Table 2: Comparison to video CoT methods. Since the models from Hu et al. [17] and Fei et al. [9] are not publicly available, we compare the results they report on Video-MME and NextQA to those of our CoF models. We also report the improvement of each approach over their baseline. Both our CoF models outperform the existing approaches on the two benchmarks.
@@ 5.2 Chain-of-frames vs chain-of-thoughts variants
+ CoT Prompting 33.5 54.7 71.5 77.0 67.4
+ SFT with QA only 31.8 54.5 73.4 64.1 57.7
+ SFT with CoT 34.3 58.6 73.7 77.9 53.1
+ SFT with CoF (ours) 36.9 59.7 76.1 79.2 71.2
@@
  • Original: InternVL2.5-4B with default prompting.
  • Original + CoT Prompting: the InternVL2.5-4B model with a prompt that encourages the model to perform intermediate reasoning before answering the question (see prompt in App. A). This approach is similar to the standard CoT used in language tasks [34].
  • Original + SFT with QA only: the InternVL2.5-4B model fine-tuned on the question-answers pairs from our CoF-Data without including the reasoning traces.
  • Original + SFT with CoT: the InternVL2.5-4B model fine-tuned on CoF-Data, where reasoning traces are included but references to specific frames are removed (e.g., “In Frame 1…” is replaced with a generic “In the video…”).
  • Original + SFT with CoF: the InternVL2.5-4B model fine-tuned on CoF-Data, i.e. our proposed approach.
@@
Results. We report results for all models on the five benchmarks in Tab. 3. First, we observe that CoT prompting alone already improves the accuracy of the original model compared to standard prompting. This result highlights the value of encouraging explicit reasoning, which video LLMs are capable of performing. Second, fine-tuning on question-answer pairs without reasoning (SFT with QA only) gives mix results, possibly due to overfitting on the training data, which might degraded the reasoning ability of the original model. Training on reasoning traces without temporal grounding (SFT with CoT) improves the results across nearly all benchmarks, with the notable exception of EventHallusion. Finally, our proposed model, CoF-InternVL2.5-4B, fine-tuned on the CoF reasoning traces, achieves the highest accuracy across all benchmarks, with improvements ranging from 4.8% to 8.7% over the original InternVL2.5-4B. These results demonstrate that our explicit temporal ground approach, chain-of-frames, offers consistent benefits across diverse tasks and domains and can serve as a strong alternative to key-frame-retrieval methods.
5.3 CoF reasoning at inference time
@@
To better understand how the CoF models utilize frame references during inference, we track the number of frames referenced in the reasoning trace per answer. In Fig. 7, we report how many answers, generated by our CoF-InternVL3-8B model across all evaluation benchmarks, contain $N$ frames for $N=1,\ldots,10$. We can see that our model learns to produce, from a relatively small training set, reasoning traces which refer to frame IDs (see qualitative examples in App. C). Moreover, the model can use frame references selectively rather than uniformly across tasks (distribution by benchmark in App. B), depending on the type of questions.
@@
Figure 7: CoF reasoning at inference time. We show the distribution of the number of frame references generated by CoF-InternVL3-8B at inference time.
@@ 6 Conclusion
We have introduced chain-of-frames (CoF), a new approach to encourage video LLMs to produce temporally grounded reasoning before providing the final answer. Compared to existing works, CoF does not require complex ad-hoc inference frameworks or auxiliary models, and we show that its training data can be extracted more efficiently and accurately from both real and synthetic videos. Our models fine-tuned on CoF-Data outperform across multiple benchmarks those obtained with alternative methods for reasoning, and even achieve results better than or similar to leading video LLMs. Overall, these features make CoF a viable option to further improve the reasoning capabilities of video LLMs. Exploring the effect of increasing the size and diversity of the training data, as well as the scale of the models, represents an exciting direction for future work.
Limitations. Our CoF training data is suitable for finetuning video LLMs that operate with a fixed FPS, ensuring that the frame indices referenced in reasoning traces align with the video frames provided to the language model. Consequently, an open question is how to adapt our method to video LLMs—such as Qwen2.5-VL [4]—that employ dynamic frame rates, while preprocessing the video.
@@ A.1 Chain-of-Frames training data
CoF from real videos (CoF-Datareal). To generate question-reasoning-answer triplets, we prompt Llama-3.1-8B-Instruct [27] using the instruction shown in Fig. 9 along with frame-aware video captions from the VideoEspresso dataset (see Fig. 3 for details). Notably, the raw video content is not included in this process. Two examples from CoF-Datareal are shown in Fig. 13.
CoF from synthetic videos (CoF-Datasynth).
The second portion of our training dataset is derived from the Clevrer dataset, which includes detailed attributes for each object in every video frame. Specifically, given a frame ID and object ID, the inside_camera_key field indicates whether the object is visible in the frame, enabling us to determine when an object enters or exits the scene. The velocity attribute reflects whether an object is moving or stationary, while the location attribute provides its absolute or relative position, which can be leveraged to estimate distances or identify collisions. The final CoF-Datasynth dataset comprises three categories of questions: Object Count, appearance order, and relative distance. Within the object count category, we define three subtypes: (i) collision-based (“How many collisions…”), (ii) motion state (“How many moving objects…”), and (iii) temporal-based, where questions reference specific segments of the video (“After object A enters…”). The questions, answers and reasoning traces are generated with the manually written templates shown in Fig. 10, making the data collection process particularly simple and fast. Examples from CoF-Datasynth are shown in Fig. 14.
Final dataset (CoF-Data).
Our training data, CoF-Data, has a total of 164,186 samples, comprising 103,683 samples from the CoF-Datareal dataset, which is based on real-world videos, and 60,503 samples from the CoF-Datasynth dataset of synthetic videos. Fig. 8 shows the distribution of how many frames are referenced in the reasoning traces, both for the final dataset and the individual splits. The CoF-Datasynth exhibits a more balanced distribution compared to the automatically generated COF-CoF-Datareal: this highlights that using synthetic videos allows us to better control various aspects of the data.